Use the function given by (f(x)=x/8+1 & g(x)=x^3 to find the indicated value (f o g)-1(5)
Answer by ewatrrr(24785)


Hi,
(f(x)=x/8+1 & g(x)=x^3
finding the indicated value (f o g)-1(5)
Step 1 substituting x^3 for x in the f(x) function
(f o g)= x^3/8 + 1
Step 2 Evaluating (f o g)(-1/5)
(f o g)(-1/5)= -.2^3/8 + 1 = -.001 + 1 = .999